# Uniqueness of asymptotically cylindrical steady gradient Ricci solitons

### v1

@@ 5.2. Roundness near infinity implies global roundness
Let $(M^{n},g,f)$ be an asymptotically $\mathbb{RP}^{n-1}$-cylindrical steady GRS of dimension $n\geq 3$. Then $(\mathbb{RP}^{n-1},\bar{g}_{\infty})$ (where $\bar{g}_{\infty}$ is from Lemma 2.6) has $\frac{n(n-1)}{2}$ linearly independent Killing fields $\{\bar{U}_{i}\}_{i=1}^{n(n-1)/2}$. Exactly as in the proof of Theorem 1.4 (see §5.1), we deduce that each far enough level set of $f$ has constant sectional curvature. The desired contradiction follows from Proposition 5.2. ∎
@@ 6. Uniqueness of Appleton’s solitons
In this section we prove Theorem 1.8.
@@ 6.1. Two basic facts
We first state two general facts that will be repeatedly used. For groups $H\subseteq G$, let $C_{G}(H)$ and $N_{G}(H)$ be the centralizer and normalizer of $H$ in $G$ respectively. As with $\mathcal{I}$ and $\mathcal{I}^{0}$, we denote identity components by a superscript 00, e.g. $C_{G}^{0}(H)$, $N_{G}^{0}(H)$.
@@
Proof.
Since $C_{G}(H)\subseteq N_{G}(H)$, we have $C_{G}^{0}(H)\subseteq N_{G}^{0}(H)$. Conversely, let $g\in N_{G}^{0}(H)$. There is a smooth curve $\gamma:[0,1]\to N_{G}^{0}(H)$ such that $\gamma(0)$ is the identity and $\gamma(1)=g$. Since $\gamma(t)\in N_{G}^{0}(H)$, we have
(6.1) $\gamma(t)h\gamma(t)^{-1}\in H\quad\text{for all }t\in[0,1]\text{ and }h\in H.$
For each fixed $h\in H$, the mapping $t\mapsto\gamma(t)h\gamma(t)^{-1}$ is smooth, and sends 00 to $h$. Since $H$ is discrete, it follows that $\gamma(t)h\gamma(t)^{-1}=h$ for all $t\in[0,1]$. Hence, $\gamma(t)\in C_{G}(H)$ for each $t\in[0,1]$. In particular, $g=\gamma(1)\in C_{G}(H)$. This implies that $N_{G}^{0}(H)\subseteq C_{G}(H)$, which further implies $N_{G}^{0}(H)\subseteq C_{G}^{0}(H)$. ∎
The next lemma is standard:
@@
Let $(\tilde{M},\tilde{g})$ be a Riemannian manifold and $\Gamma\subset\mathcal{I}(\tilde{M},\tilde{g})$ a finite subgroup of isometries acting freely. Then the quotient manifold $(M,g)$ has isometry group
@@ 6.2. Isometries of Berger metrics
6.3. Proof of Theorem 1.8
The basic idea is similar to the proofs of Theorems 1.4 and 1.6, but rather than showing that the level sets of $f$ have constant sectional curvature, we can only show they are isometric to $(L_{m,k},g_{a,b})$ using Corollary 6.7. We will combine this with the next few results to obtain the desired conclusions.
@@
Let $(M,g,f)$ be an asymptotically $L_{m,k}$-cylindrical steady GRS with $m\geq 1$ and $k\geq 3$. Then $(L_{m,k},\bar{g}_{\infty})$ (where $\bar{g}_{\infty}$ is from Lemma 2.6) has $(m+1)^{2}$ linearly independent Killing fields $\{\bar{U}_{i}\}_{i=1}^{(m+1)^{2}}$ by Lemma 6.4. Since $(L_{m,k},\bar{g}_{\infty})$ is homogeneous, the Killing fields span the tangent space of $L_{m,k}$ at each point.
Thus, Corollary 3.2 gives $(m+1)^{2}$ linearly independent Killing fields $\{W_{i}\}_{i=1}^{(m+1)^{2}}$ on $\{f=\rho\}$ for all large $\rho$, which span the tangent space to $\{f=\rho\}$ at every point. Hence, all far enough level sets of $f$ are diffeomorphic to $L_{m,k}$ and are equipped with a homogeneous metric with isometry group of dimension at least $(m+1)^{2}$. By Corollary 6.7, each of these level sets is isometric to some $(L_{m,k},g_{a,b})$. Therefore:
@@
Thus, we may apply Proposition 4.2 with $G=U(m+1)$ and $H=U(m)\times\mathbb{Z}_{k}$ to get the following.
@@
It remains to examine the boundary conditions. By (iii) and Corollary 6.9, we have either $K=U(m)\times\mathbb{Z}_{2k}$ (giving $X\cong L_{m,2k}$), or $K=U(m)\times U(1)$ (giving $X\cong\mathbb{CP}^{m}$). However, the first option is incompatible with (iv) because it is impossible for $(G/H,g_{a(s),b(s)})$ to converge in the Hausdorff sense to a metric space homeomorphic to $L_{m,2k}$. Hence $K=U(m)\times U(1)$ and $X\cong\mathbb{CP}^{m}$.
@@ 7.1. Homogeneous metrics on $SU(2)/\Gamma$
We now prove the non-existence of asymptotically $SU(2)/\Gamma$-cylindrical steady GRSs for $\Gamma\in\{\tilde{T},\tilde{O},\tilde{I}\}$.
@@
Let $(M^{n},g,f)$ be an asymptotically $SU(2)/\Gamma$-cylindrical steady GRS, where $\Gamma\in\{\tilde{T},\tilde{O},\tilde{I}\}$. By Corollary 7.2, $(SU(2)/\Gamma,\bar{g}_{\infty})$ (where $\bar{g}_{\infty}$ is from Lemma 2.6) has 3 linearly independent Killing fields $\{\bar{U}_{1},\bar{U}_{2},\bar{U}_{3}\}$ spanning the tangent space at each point. Thus, Corollary 3.2 gives 3 linearly independent Killing fields $\{W_{1},W_{2},W_{3}\}$ on $\{f=\rho\}$ for all large $\rho$, which span the tangent space to $\{f=\rho\}$ at every point. Hence, all far enough level sets of $f$ are diffeomorphic to $SU(2)/\Gamma$ and are equipped with homogeneous metrics. By Lemma 7.3, the level sets have constant sectional curvature. The desired contradiction follows from Proposition 5.2. ∎
7.3. Uniqueness of $\mathbb{Z}_{2}$-quotients of Appleton’s solitons
Finally, we turn to asymptotically $SU(2)/\tilde{D}_{k}$-cylindrical steady GRSs for $k\geq 2$. There turn out to exist such GRSs, given by $\mathbb{Z}_{2}$-quotients of Appleton solitons in $\mathcal{A}_{1,2k}$. To see this, we use the notation of §1.2. For any $a,b>0$, the manifold $(L_{1,2k},g_{a,b})$ is the quotient of $(SU(2),\tilde{g}_{a,b})$ by the right action of the subgroup $\mathbb{Z}_{2k}$ generated by $e^{\pi i/k}$. By Lemma 7.1, right multiplication by $j$ belongs to $\mathcal{I}(SU(2),\tilde{g}_{a,b})$. This descends to a fixed-point-free isometric involution of $(L_{1,2k},g_{a,b})$. Hence, the quotient of $(L_{1,2k},g_{a,b})$ by this involution is exactly the quotient of $(SU(2),\tilde{g}_{a,b})$ by the right action of $\tilde{D}_{k}$. Denote this quotient by $(SU(2)/\tilde{D}_{k},\hat{g}_{a,b})$.
@@ 7.3. Uniqueness of $\mathbb{Z}_{2}$-quotients of Appleton’s solitons
Now, we prove the converse: every asymptotically $SU(2)/\tilde{D}_{k}$-cylindrical steady GRS must be of the form $A/\mathbb{Z}_{2}$.
@@
Let $(M,g,f)$ be an asymptotically $SU(2)/\tilde{D}_{k}$-cylindrical steady GRS with $k\geq 2$. Following the proof of Theorem 1.10, we deduce that all far enough level sets of $f$ are diffeomorphic to $SU(2)/\tilde{D}_{k}$ and are equipped with homogeneous metrics. Therefore:
@@
Thus, we may apply Proposition 4.2 with $G=SU(2)$ and $H=\tilde{D}_{k}$ to get the following.
@@
    There is a closed subgroup $K$ of $SU(2)$ containing $H$ such that $K/H\cong\mathbb{S}^{\ell}$ for some $\ell\geq 0$, and $X\cong SU(2)/K$.
@@
By (iii) and the classification of closed subgroups of $SU(2)$ (see e.g. [cmp-su2]), we have $K=\tilde{D}_{2k}$ (giving $X\cong SU(2)/\tilde{D}_{2k}$) or $K=\mathrm{Pin}(2)$ (giving $X\cong\mathbb{RP}^{2}$). However, since the covering metrics of $g(s)$ are given by (7.16), the Hausdorff limit of $(SU(2)/\tilde{D}_{k},g(s))$ as $s\to 0$ must be homeomorphic to either $SU(2)/\tilde{D}_{k}$ or a space of lower dimension. This and (iv) rule out $K=\tilde{D}_{2k}$. Thus, $K=\mathrm{Pin}(2)$ and $X\cong\mathbb{RP}^{2}$. Then $M$ is diffeomorphic to the total space of a vector bundle over $\mathbb{RP}^{2}$ whose sphere bundles are diffeomorphic to $SU(2)/\tilde{D}_{k}$. This is exactly the non-orientable 2-plane bundle over $\mathbb{RP}^{2}$ with Euler number ±$\pm k$. We have $\pi_{1}(M)=\pi_{1}(\mathbb{RP}^{2})=\mathbb{Z}_{2}$.
@@
A smooth double cover of $M$ is obtained by pulling back the bundle $M$ by the antipodal map $\mathbb{S}^{2}\to\mathbb{RP}^{2}$. This is a 2-plane bundle $\tilde{M}$ over $\mathbb{S}^{2}$ with Euler number ±$\pm 2k$, i.e. the total space of $\mathcal{O}_{\mathbb{CP}^{1}}(-2k)$, whose sphere bundles are diffeomorphic to $L_{1,2k}$. Lifting the soliton structure from $M$ to $\tilde{M}\cong\mathcal{O}_{\mathbb{CP}^{1}}(-2k)$ and using (ii), we see that the lifted soliton $(\tilde{M},\tilde{g},\tilde{f})$ is an asymptotically $L_{1,2k}$-cylindrical steady GRS. By Theorem 1.8, we have $(\tilde{M},\tilde{g},\tilde{f})\in\mathcal{A}_{1,2k}$. ∎
Appendix A Asymptotic cylindricity of Appleton’s solitons
In this appendix we prove:
Proposition A.1.
Let $m,k\geq 1$. Then any $(M,g,f)\in\mathcal{A}_{m,k}$ is asymptotically $L_{m,k}$-cylindrical.
Proving this is a matter of assembling various facts from Appleton’s analysis in [appleton]; there it was already shown that any $(M,g,f)\in\mathcal{A}_{m,k}$ has the same asymptotic profile as the Bryant soliton quotiented by $\mathbb{Z}_{k}$. It is more convenient to use the following equivalent definition of asymptotic cylindricity:
Definition A.2.
Let $\mathcal{S}$ be a spherical space form of dimension $n-1$. A steady GRS $(M^{n},g,f)$ is asymptotically $\mathcal{S}$-cylindrical if for any sequence of points $p_{i}\in M$ going to infinity, we have convergence of pointed manifolds
(A.1) $(M,R(p_{i})g,p_{i})\to(\mathcal{S}\times\mathbb{R},\bar{g}_{\mathcal{S}}+dz\otimes dz,p_{\infty})$
in the Cheeger–Gromov sense, without passing to a subsequence. Here, $R$ is the scalar curvature of $g$, $\bar{g}_{\mathcal{S}}$ is a metric on $\mathcal{S}$ with constant sectional curvature normalized to have scalar curvature 1, and $p_{\infty}$ is a point on $\mathcal{S}\times\mathbb{R}$.
Let $m,k\geq 1$, and fix $(M,g,f)\in\mathcal{A}_{m,k}$. Recall from §1.2 that outside the $\mathbb{CP}^{m}$ core, $(M,g)$ is $(0,\infty)\times L_{m,k}$ equipped with a metric of the form
(A.2) $g=ds^{2}+g_{a(s),b(s)},$
where $g_{a(s),b(s)}$ is the $\mathbb{Z}_{k}$-quotient of the Berger metric $\tilde{g}_{a(s),b(s)}$ on $\mathbb{S}^{2m+1}$. Let $Q=a/b$.
Lemma A.3 ([appleton]).
The functions $a$ and $b$ satisfy the ODEs
(A.3) $a^{\prime\prime}$ $=\frac{2mQ^{4}}{a}-2m\frac{(a^{\prime})^{2}}{a}-\left(f^{\prime}-2m\frac{Q^{\prime}}{Q}\right)a^{\prime},$
(A.4) $b^{\prime\prime}$ $=2\frac{m+1-Q^{2}}{b}-2m\frac{(b^{\prime})^{2}}{b}-\left(f^{\prime}+\frac{Q^{\prime}}{Q}\right)b^{\prime}.$
Moreover, after possibly rescaling the soliton $(M,g,f)$, it holds that
(A.5) $\lim_{s\to\infty}Q$ $=1,$
(A.6) $\lim_{s\to\infty}Q^{\prime}$ $=0,$
(A.7) $\lim_{s\to\infty}f^{\prime}$ $=1,$
and
(A.8) $a(s)\sim\sqrt{4ms},\quad b(s)\sim\sqrt{4ms}\quad\text{as }s\to\infty.$
Proof.
The ODEs (A.3) and (A.4) follow from (1.8) and (1.9) respectively, while (A.5) holds by the definition of $\mathcal{A}_{m,k}$. The limit (A.6) is due to [appleton]*Lemma 6.4. The limit $\lim_{s\to\infty}f^{\prime}$ exists and is positive by [appleton]*Lemma 4.2. By scaling $g$, it can be arranged to equal $1$, establishing (A.7). Finally, (A.8) is [appleton]*Corollary 6.7. ∎
The next ODE lemma follows from the proof of [appleton]*Lemma 6.2:
Lemma A.4.
Let $y:[0,\infty)\to\mathbb{R}$ be a solution to the ODE
(A.9) $y^{\prime\prime}=\frac{c_{1}(s)}{2y}-\alpha\frac{(y^{\prime})^{2}}{y}-c_{2}(s)y^{\prime},$
where $\alpha>1$ is a constant, and $c_{i}:[0,\infty)\to\mathbb{R}$, $i=1,2$, are two positive functions converging to positive limits $c_{i}^{*}>0$ as $s\to\infty$. Then the function $z=y^{\alpha+1}$ satisfies
(A.10) $\frac{z^{\prime}}{z^{\frac{\alpha-1}{\alpha+1}}}\sim\frac{\alpha+1}{2}\frac{c_{1}^{*}}{c_{2}^{*}}\quad\text{as }s\to\infty.$
Proof of Proposition A.1.
Let $(M,g,f)\in\mathcal{A}_{m,k}$ with $m,k\geq 1$. Using Lemma A.3, we can apply Lemma A.4 to $a$ and $b$, giving
(A.11) $aa^{\prime}\sim 2m,\quad bb^{\prime}\sim 2m\quad\text{as }s\to\infty.$
Then by (A.8),
(A.12) $a^{\prime}(s)\sim\frac{1}{\sqrt{ms}},\quad b^{\prime}(s)\sim\frac{1}{\sqrt{ms}}\quad\text{as }s\to\infty.$
Inserting this into (A.3) and (A.4), and also using the other limits from Lemma A.3, we find that the leading terms cancel and hence
(A.13) $a^{\prime\prime}(s)=o(s^{-\frac{1}{2}}),\quad b^{\prime\prime}(s)=o(s^{-\frac{1}{2}}).$
Next, we use the formula
(A.14) $R=-2m\frac{a^{\prime\prime}}{a}-4m\frac{b^{\prime\prime}}{b}-4m\frac{a^{\prime}b^{\prime}}{ab}-2m(2m-1)\left(\frac{b^{\prime}}{b}\right)^{2}-2m\frac{a^{2}}{b^{4}}+\frac{2m(2m+2)}{b^{2}}$
(see [appleton]*Appendix A) and the above asymptotics to see that
(A.15) $R\sim\frac{2m+1}{2s}.$
By repeatedly differentiating (A.3) and (A.4), one can obtain for each $k\in\mathbb{N}$
(A.16) $a^{(k)}(s)=o(1),\quad b^{(k)}(s)=o(1).$
For each $i\in\mathbb{N}$, define the functions $a_{i}:\mathbb{R}\to\mathbb{R}$ and $b_{i}:\mathbb{R}\to\mathbb{R}$ by
(A.17) $a_{i}(z)$ $=\sqrt{R(s_{i})}a\left(s_{i}+\frac{z}{\sqrt{R(s_{i})}}\right),$
(A.18) $b_{i}(z)$ $=\sqrt{R(s_{i})}b\left(s_{i}+\frac{z}{\sqrt{R(s_{i})}}\right).$
By (A.15) and (A.16), it holds for each $k\geq 0$ that
(A.19) $a_{i}\to\sqrt{2m(2m+1)},\quad b_{i}\to\sqrt{2m(2m+1)}\quad\text{in }C^{k}_{\mathrm{loc}}(\mathbb{R})$
as $i\to\infty$.
Now, let $p_{i}\in M$ be a sequence of points such that $s_{i}=s(p_{i})\to\infty$. Since the scalar curvature (A.14) is a function of $s$, we have $R(p_{i})=R(s_{i})$. Then for each $i$,
(A.20) $R(p_{i})g$ $=R(s_{i})(ds^{2}+g_{a(s),b(s)})$
(A.21) $=dz^{2}+R(s_{i})g_{a\left(s_{i}+\frac{z}{\sqrt{R(s_{i})}}\right),b\left(s_{i}+\frac{z}{\sqrt{R(s_{i})}}\right)}$
(A.22) $=dz^{2}+g_{a_{i}(z),b_{i}(z)},$
where $z=\sqrt{R(s_{i})}(s-s_{i})$.
The metric on $L_{m,k}$ given by
(A.23) $\bar{g}=g_{\sqrt{2m(2m+1)},\sqrt{2m(2m+1)}}$
has constant sectional curvature, and is normalized to have constant scalar curvature 1. Thus, to obtain the required pointed convergence (A.1), it suffices to check that for all $k\geq 0$ and $\rho>0$, the sequence
(A.24) $\sup_{|z|\leq\rho}\left\lVert g_{a_{i}(z),b_{i}(z)}-\bar{g}\right\rVert_{C^{k}(\bar{g})}$
converges to 00 as $i\to\infty$. This convergence indeed holds because of (A.19). ∎

### v2

@@ 5.3. Uniqueness of the $\mathbb{Z}_{2}$-quotient of the Bryant soliton
Let $(M^{n},g,f)$ be an asymptotically $\mathbb{RP}^{n-1}$-cylindrical steady GRS on a smooth manifold of dimension $n\geq 3$. Then $(\mathbb{RP}^{n-1},\bar{g}_{\infty})$ (where $\bar{g}_{\infty}$ is from Lemma 2.6) has $\frac{n(n-1)}{2}$ linearly independent Killing fields $\{\bar{U}_{i}\}_{i=1}^{n(n-1)/2}$. As in the proof of Theorem 1.4 (see §5.1), we deduce that each far enough level set of $f$ has constant sectional curvature. Proposition 5.2 now gives a contradiction. ∎
The same argument also proves part (b) of Corollary 1.10.
@@ 6. Uniqueness of Appleton’s solitons
In this section we prove Theorem 1.8, as well as part (c) of Corollary 1.10.
@@ 6.1. Two basic facts
For groups $H\subseteq G$, let $C_{G}(H)$ and $N_{G}(H)$ be the centralizer and normalizer of $H$ in $G$ respectively. As with $\mathcal{I}$ and $\mathcal{I}^{0}$, we denote identity components by a superscript 00, e.g. $C_{G}^{0}(H)$, $N_{G}^{0}(H)$. We first state two general facts that will be repeatedly used. Their proofs are rudimentary and therefore omitted.
@@
The following is a standard fact about isometry groups of quotient manifolds:
@@
Let $(\tilde{M},\tilde{g})$ be a simply-connected Riemannian manifold and $\Gamma\subset\mathcal{I}(\tilde{M},\tilde{g})$ a finite subgroup of isometries acting freely. Then the quotient manifold $(M,g)$ has isometry group
@@ 6.2. Isometries of Berger metrics
6.3. Uniqueness of Appleton’s solitons
The proof of Theorem 1.8 is similar to those of Theorems 1.4 and 1.6, but rather than showing that the level sets of $f$ have constant curvature, we can only show they are isometric to $(L_{m,k},g_{a,b})$ using Corollary 6.7. We will combine this with the next few results to obtain the desired conclusions.
@@ 6.3. Uniqueness of Appleton’s solitons
Let $(M,g,f)$ be an asymptotically $L_{m,k}$-cylindrical steady GRS on a smooth manifold, where $m\geq 1$ and $k\geq 3$. Then $(L_{m,k},\bar{g}_{\infty})$ (where $\bar{g}_{\infty}$ is from Lemma 2.6) has $(m+1)^{2}$ linearly independent Killing fields $\{\bar{U}_{i}\}_{i=1}^{(m+1)^{2}}$ by Lemma 6.4. Since $(L_{m,k},\bar{g}_{\infty})$ is homogeneous, the Killing fields span the tangent space of $L_{m,k}$ at each point.
Corollary 3.2 gives $(m+1)^{2}$ linearly independent Killing fields for all large $\rho$, which span the tangent space to $\{f=\rho\}$ at every point. Hence, all far enough level sets of $f$ are diffeomorphic to $L_{m,k}$ and are equipped with a homogeneous metric with isometry group of dimension at least $(m+1)^{2}$. By Corollary 6.7, each of these level sets is isometric to some $(L_{m,k},g_{a,b})$. Therefore:
@@
Using Proposition 4.2 with $G=U(m+1)$ and $H=U(m)\times\mathbb{Z}_{k}$, we get the following.
@@
It remains to examine the boundary conditions. By (iii) and Corollary 6.9, there are two possibilities for $K$ and $X$:
  1. (a)
    $K=U(m)\times\mathbb{Z}_{2k}$ and $X\cong L_{m,2k}$.
  2. (b)
    $K=U(m)\times U(1)$ and $X\cong\mathbb{CP}^{m}$.
However, the first option is incompatible with (iv) because it is impossible for $(G/H,g_{a(s),b(s)})$ to converge in the Hausdorff sense to a metric space homeomorphic to $L_{m,2k}$. Hence $K=U(m)\times U(1)$ and $X\cong\mathbb{CP}^{m}$.
@@
Proof of part (c) of Corollary 1.10.
Let $(M^{4},g,f)$ be an asymptotically $L_{1,k}$-cylindrical steady GRS singularity model on a smooth orbifold with isolated singularities, where $k\geq 3$. The proof of Theorem 1.8 applies verbatim, though in item (iii) we also allow $K/H$ to be a spherical space form $\mathcal{S}^{\prime}$ of dimension $\ell\geq 0$. Using Lemma 6.8, we see that the only new possibility for $(K,X)$ is that $K=U(2)$ and $X$ is a point. Then $M$ is diffeomorphic to a cone over $L_{1,k}$. We now run the same argument used in the last paragraph of the proof of Proposition 5.2. This shows that $(M,g,f)$ is the Bryant soliton quotiented by $\mathbb{Z}_{k}$. ∎
@@ 7.1. Homogeneous metrics on $SU(2)/\Gamma$
We now prove the uniqueness of asymptotically $SU(2)/\Gamma$-cylindrical steady GRSs for $\Gamma\in\{\tilde{T},\tilde{O},\tilde{I}\}$.
@@
Let $(M^{n},g,f)$ be an asymptotically $SU(2)/\Gamma$-cylindrical steady GRS on a smooth orbifold with isolated singularities, where $\Gamma\in\{\tilde{T},\tilde{O},\tilde{I}\}$. By Corollary 7.2, $(SU(2)/\Gamma,\bar{g}_{\infty})$ (where $\bar{g}_{\infty}$ is from Lemma 2.6) has 3 linearly independent Killing fields $\{\bar{U}_{1},\bar{U}_{2},\bar{U}_{3}\}$ spanning the tangent space at each point. Thus, Corollary 3.2 gives 3 linearly independent Killing fields on $\{f=\rho\}$ for all large $\rho$, which span the tangent space to $\{f=\rho\}$ at every point. Hence, all far enough level sets of $f$ are diffeomorphic to $SU(2)/\Gamma$ and are equipped with homogeneous metrics. By Lemma 7.3, the level sets have constant sectional curvature. The result follows from Proposition 5.2. ∎
7.3. $\mathbb{Z}_{2}$-quotients of Appleton’s solitons
Finally, we turn to asymptotically $SU(2)/\tilde{D}_{k}$-cylindrical steady GRSs for $k\geq 2$. There turn out to exist such GRSs on smooth manifolds, given by $\mathbb{Z}_{2}$-quotients of Appleton solitons in $\mathcal{A}_{1,2k}$. To see this, we use the notation of §1.2. For any $a,b>0$, the manifold $(L_{1,2k},g_{a,b})$ is the quotient of $(SU(2),\tilde{g}_{a,b})$ by the right action generated by $e^{\pi i/k}$. By Lemma 7.1, right multiplication by $j$ belongs to $\mathcal{I}(SU(2),\tilde{g}_{a,b})$. This descends to a fixed-point-free isometric involution of $(L_{1,2k},g_{a,b})$. Hence, the quotient of $(L_{1,2k},g_{a,b})$ by this involution is exactly the quotient of $(SU(2),\tilde{g}_{a,b})$ by the right action of $\tilde{D}_{k}$. Denote this quotient by $(SU(2)/\tilde{D}_{k},\hat{g}_{a,b})$.
@@ 7.3. $\mathbb{Z}_{2}$-quotients of Appleton’s solitons
7.4. Proof of Theorem 1.12
We prove that $A/\mathbb{Z}_{2}$ are the only asymptotically $SU(2)/\tilde{D}_{k}$-cylindrical steady GRSs in addition to Bryant soliton quotients.
@@
Let $(M,g,f)$ be an asymptotically $SU(2)/\tilde{D}_{k}$-cylindrical steady GRS on a smooth orbifold with isolated singularities, where $k\geq 2$. Following the proof of Theorem 1.11, we deduce that all far enough level sets of $f$ are diffeomorphic to $SU(2)/\tilde{D}_{k}$ and are equipped with homogeneous metrics. Therefore:
@@
Then applying Proposition 4.2 with $G=SU(2)$ and $H=\tilde{D}_{k}$ gives the following:
@@
    There is a closed subgroup $K$ of $SU(2)$ containing $H$ such that $K/H\cong\mathcal{S}^{\prime}$ for a spherical space form $\mathcal{S}^{\prime}$ of dimension $\ell\geq 0$. We have $X\cong SU(2)/K$.
@@
By (iii) and the classification of closed subgroups of $SU(2)$ (see e.g. [cmp-su2]), the possibilities for $K$ and $X$ are as follows:
  1. (a)
    $K=\tilde{D}_{2k}$ and $X\cong SU(2)/\tilde{D}_{2k}$.
  2. (b)
    $K=\mathrm{Pin}(2)$ and $X\cong\mathbb{RP}^{2}$.
  3. (c)
    $K=SU(2)$ and $X$ is a point.
However, since the covering metrics of $g(s)$ are given by (7.16), the Hausdorff limit of $(SU(2)/\tilde{D}_{k},g(s))$ as $s\to 0$ is homeomorphic to either $SU(2)/\tilde{D}_{k}$ or a space of lower dimension. This and (iv) rule out $K=\tilde{D}_{2k}$, so we consider only the other two cases:
@@
Case (b): $K=\mathrm{Pin}(2)$ and $X\cong\mathbb{RP}^{2}$. In this case $M$ is diffeomorphic to the total space of a vector bundle over $\mathbb{RP}^{2}$ whose sphere bundles are diffeomorphic to $SU(2)/\tilde{D}_{k}$. This is exactly the non-orientable 2-plane bundle over $\mathbb{RP}^{2}$ with Euler number ±$\pm k$. In particular, $M$ is a manifold with $\pi_{1}(M)=\pi_{1}(\mathbb{RP}^{2})=\mathbb{Z}_{2}$.
@@
A smooth double cover of $M$ is obtained by pulling back the bundle $M$ by the antipodal map $\mathbb{S}^{2}\to\mathbb{RP}^{2}$. This is a 2-plane bundle $\tilde{M}$ over $\mathbb{S}^{2}$ with Euler number ±$\pm 2k$, i.e. the total space of $\mathcal{O}_{\mathbb{CP}^{1}}(-2k)$, whose sphere bundles are diffeomorphic to $L_{1,2k}$. Lifting the soliton structure from $M$ to $\tilde{M}\cong\mathcal{O}_{\mathbb{CP}^{1}}(-2k)$ and using (ii), we see that the lifted soliton $(\tilde{M},\tilde{g},\tilde{f})$ is an asymptotically $L_{1,2k}$-cylindrical steady GRS. By Theorem 1.8, we have $(\tilde{M},\tilde{g},\tilde{f})\in\mathcal{A}_{1,2k}$.
Case (c): $K=SU(2)$ and $X$ is a point. In this case, $M$ is diffeomorphic to a cone over $\mathcal{S}$. We now run the same argument used in the last paragraph of the proof of Proposition 5.2. This shows that $(M,g,f)$ is the Bryant soliton quotiented by $\tilde{D}_{k}$. ∎